Which equation is NOT a direct variation?
A) y = 50x
B) 5x + 2y= 10
C) - 2y = x
D) - 3x + 2y = 0

Answer by Theo(13342)

a direct variation is of the form y = kx.

k is the constant of variation.

y = 50x is clearly a direct variation with k = 50.

2y = x looks like a direct variation because, when you solve for y, you get y = 1/2 * x, where k is equal to 1/2, so it fits the form of y = kx.

-3x + 2y = 0 leads to y = 3/2 * x, so it also fits the form of y = kx where k = 3/2.

that leaves 5x + 2y = 10 which has to be the answer.

solve for y to get y = (10 - 5x) / 2 which is equal to y = 5 - 5/2 * x which does not fit the form of y = kx.

so, i would say selection B is the one that is not a direct variation, because it doesn't fit the general form of a direct variation equation.